OpenStudy (meilendurcer):

Rosa, Roberto, Andrea, and Inno find an estimate for . Who has proposed the best solution?

OpenStudy (meilendurcer):

\[\sqrt{10}\]

OpenStudy (anonymous):

What are their estimates?

OpenStudy (meilendurcer):

Rosa: "Use\[\sqrt{9}\] and \[\sqrt{25}\] to estimate." Roberto: "I will use \[\sqrt{4}\] and \[\sqrt{9}\]." Andrea: "It should be \[\sqrt{11}\] and \[\sqrt{12}\]." Inno: "Use \[\sqrt{9}\] and \[\sqrt{16}\] to find the estimate

OpenStudy (meilendurcer):

sorry it's long

OpenStudy (stamp):

@meilendurcer What do YOU think the answer is? Explain your answer. ("I don't know" is not an acceptable response)

OpenStudy (meilendurcer):

to be truthful i was gonna say Roberto

OpenStudy (stamp):

sqrt 4 and sqrt of 9 would be a good method to estimate the sqrt of 5 through 8 Because think, you can evaluate the sqrt of 4 and sqrt of 9, but not sqrt of 5. So you could say: Since 5 is inbetween 4 and 9, the square root of 5 is between the square root of 4 and the square root of 9 \[\sqrt4,\ \sqrt 5,\ \sqrt9\]simplify\[2,\ \sqrt5,\ 3\]So the square root of 5 is somewhere between the values 2 and 3, probably closer to the 2 side.

OpenStudy (stamp):

Notice 10 is not between 4 and 9, so those are not helpful for us if we want to estimate the square root of 10 relative to the square root of 4 and square root of 9.

OpenStudy (meilendurcer):

okay second choice Inno?

OpenStudy (stamp):

@meilendurcer What would you estimate sqrt 10 's value to be

OpenStudy (stamp):

Arrange\[\sqrt9,\sqrt10,\sqrt16\]Simplify\[3,\sqrt10,\ 4\]The term square root of ten is somewhere between 3 and 4, more likely closer to the three value.

OpenStudy (stamp):

Lo and behold, a calculator returns the value of: http://www.wolframalpha.com/input/?i=square+root+of+10 ~ 3.16 So you see, we properly estimated the value. I hope you know understand.